OpenStudy (anonymous):

F(x) = (5x^3 + 4)^3 find second deravitive F’(x) = 3(5x^3 + 4)3-1 (15x^2) = (45x2) (5x^3 + 4)2 F”(X) = (45x^2) (5x^3 + 4)2 (45x^2)*2(5x^3 + 4)*15x2] + [90x * 90(5x^3 + 4)^2] how to open this up

OpenStudy (amistre64):

my eyes are burning!!

OpenStudy (amistre64):

F(x) = (5x^3 + 4)^3 F'' = 3(5x^3 + 4)^2 . 15x^2

OpenStudy (watchmath):

\(F(x)=(5x^3+4)^3\) \(F'(x)=3(5x^3+4)^2\cdot 15x^2\) \(F''(x)=90x\cdot (5x^3+4)^2+90x^2(5x^3+4)\cdot 15x^2\)

OpenStudy (amistre64):

that was sposed to be F'.... F'' = 3(5x^3 + 4)^2 . 30x + 6(5x^3 + 4) 15x^2 15x^2

OpenStudy (anonymous):

right now what after that step

OpenStudy (amistre64):

that is the second derivative

OpenStudy (anonymous):

how do you open that up

OpenStudy (amistre64):

you only derive twice to find the second derivative....

OpenStudy (anonymous):

understood how do you open teacher wants us to open

OpenStudy (amistre64):

if you wanna expand it; then multiply it out..... right?

OpenStudy (amistre64):

90x(25x^6 + 16 + 20x^3) + 15.15.6.x^4(5x^3 +4)

OpenStudy (amistre64):

90x(25x^6 + 16 + 20x^3) + 1350x^4(5x^3 +4) 90(25)x^7 + (90x)16 + (90)20x^4 + 1350(5)x^7 +4(1350x^4) if i did it right

OpenStudy (anonymous):

your first derivative is wrong it should be 3(5x^3 + 4)3-1 (15x^2) = (45x2) (5x^3 + 4)^2

OpenStudy (anonymous):

3(5x^3 + 4)^2 (15x^2) = (45x^2) (5x^3 + 4)^2

OpenStudy (amistre64):

F' = 3(5x^3 + 4)^2 . 15x^2 this? not according to what you gave us...

OpenStudy (anonymous):

(5x^3 + 4)^3

OpenStudy (anonymous):

original right

OpenStudy (anonymous):

derivative of inner

OpenStudy (anonymous):

* chain rule

OpenStudy (amistre64):

F' = 45x^2 (5x-4)^2 ... we agree :)

OpenStudy (anonymous):

k lol

OpenStudy (anonymous):

now 2nd derivative

OpenStudy (amistre64):

to make things easier; we could just expand this right here and take the derivative afterwards

OpenStudy (anonymous):

smart

OpenStudy (anonymous):

so

OpenStudy (amistre64):

i have my moments lol

OpenStudy (anonymous):

so 25x - 40x + 16

OpenStudy (amistre64):

45x^2(25x^6 -40x^3 +16) right?

OpenStudy (anonymous):

* 45x^2

OpenStudy (anonymous):

how to the 6th

OpenStudy (amistre64):

1125x^8 -1800x^5 + 720x^2

OpenStudy (amistre64):

5.5.x^3.x^3 = 25x^6

OpenStudy (anonymous):

its to the ^2

OpenStudy (amistre64):

-4.-4 = +16 5.-4.2.x^3 = -40x^3

OpenStudy (amistre64):

just the innards; exapnding the ^2 part....

OpenStudy (amistre64):

then multiply in the 45x^2 after that

OpenStudy (anonymous):

oh you did product + chain rule

OpenStudy (amistre64):

1125x^8 -1800x^5 + 720x^2 is what you should get

OpenStudy (anonymous):

check this link http://www.wolframalpha.com/input/?i=d^2%2Fdx^2%28%285x^3+%2B+4%29^3%29 it says diff

OpenStudy (amistre64):

45x^2 (5x^3 -4)^2 45x^2(25x^6 -40x^3 +16) 1125x^8 -1800x^5 + 720x^2 all i did was expand it to get to here; no second derivative yet..

OpenStudy (anonymous):

ahhh

OpenStudy (amistre64):

now we can easily...or easilier... take the deriavtive of this

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

1125(8)x^7 -1800(5)x^4 + 720(2)x 9000x^7 -9000x^4 + 1440x if i kept track of it all

OpenStudy (amistre64):

factor out the 360x now

OpenStudy (amistre64):

and it should match wolframs

OpenStudy (amistre64):

360x (25x^6 -25x^3 +4)

OpenStudy (amistre64):

only diff is that I got a negative from a typo in the beginning

OpenStudy (anonymous):

thanks man

OpenStudy (amistre64):

youre welcome :)